# Lossy Source Coding with Broadcast Side Information

### v1

@@ I Introduction
This paper considers the multi-user lossy compression of a source $S$ with two receivers and a helper. However, the helper who observes a correlated source $T$ can only communicate to the receivers through a noisy broadcast channel. A similar problem studied in [13] considers the case where two senders observe the same source and communicate to the receiver through two separate noisy channels. Sufficient and necessary conditions for the separation-based scheme to be optimal were derived. We extend the model by considering a pair of correlated sources and two receivers, and simplify the link from the source encoder to the receiver to a noiseless link.
When the noisy channel is a degraded broadcast channel, our model also shares some similarities with source coding with degraded side information [6, 16]. For some special cases, our model in fact reduces to source coding with degraded SI when an uncoded scheme is used. We compare these two problems with a quadratic Gaussian example. [^1] Throughout this paper, random variables, sample values and their alphabets are denoted by capital, lower and calligraphic letters, e.g., $X,x$ and $\mathcal{X}$. A sequence with length $n$ is denoted by $X^{n}=(X_{1},X_{2},...,X_{n})$.
@@
The sender $TX_{1}$ observes the source $S^{n}$ and sends a lossy description to both receivers at rate $R$. The helper observes a correlated source $T^{n}$. It encodes the source into some codeword $X^{l}$ and inputs it to a degraded broadcast channel $P_{YZ|X}=P_{Y|X}P_{Z|Y}$. The receivers observe channel output $Z^{l}$ and $Y^{l}$ and reconstruct $\hat{S}^{n}_{1}$ and $\hat{S}^{n}_{2}$ with distortion levels $D_{1}$ and $D_{2}$, respectively, such that
@@
In some of the results, we restrict the helper to use a separation-based coding scheme. That says, the helper first quantizes the source $T^{n}$ using a lossy source code, and then encodes it using a channel code. The formal definitions of the general joint-source channel coding and separation-based coding scheme are as follows.
@@
  • A pair of decoders $g_{1}:\mathcal{M}\times\mathcal{Z}^{l}\to\hat{S}^{n}_{1}$ and $g_{2}:\mathcal{M}\times\mathcal{Y}^{l}\to\hat{S}^{n}_{2}$.
@@
This section provides an outer bound of the rate-distortion-bandwidth function as well as some inner bounds with additional assumptions.
@@
The outer bound uses a remote source argument that was also used in [17, 7, 20, 12]. The bound is motivated by the rate loss in the source coding with a helper problem. A simple example is the lossless source coding with a helper. The optimal region for this model is
@@
Let $\mathcal{R}^{in}$ be the set of quadruple $(R,D_{1},D_{2},\rho)$ such that there exists a set of finite auxiliary random variables $(U_{1},U_{2},V_{1},V_{2},W)$ such that
@@
The RDB region of the lossy source coding with broadcasted side information under the separation assumption satisfies $\mathcal{R}^{in}\subseteq\mathcal{R}.$ Furthermore, the cardinalities of the auxiliary random variable alphabets satisfy
@@
The converse is given in Section VI. The achievability follows by constructing lossy descriptions $\{\hat{s}^{n}_{1}\},\{\hat{s}^{n}_{2}\}$ and dividing their indices into two parts, sent by the sender and the helper, respectively. The details are provided in Appendix B. In the following, we consider a special case in which the distortion function at the weak receiver is a deterministic distortion function[2][18].
@@
Given a function $\psi:\mathcal{S}\to\widetilde{S}$ of the source, a distortion function $d:\mathcal{S}\times\hat{\mathcal{S}}\to[0,+\infty)$ is called a deterministic distortion if

### v2

@@ I Introduction
This paper considers the multi-user lossy compression of a source $S$ with two receivers and a helper. However, the helper who observes a correlated source $T$ can only communicate to the receivers through a noisy broadcast channel. A similar problem studied in [13] considers the case where two senders observe the same source $S$ and communicate to the receiver through two separate noisy channels. Sufficient and necessary conditions for the separation-based scheme to be optimal were derived. We extend the model by considering a pair of correlated sources $(S,T)$ and two receivers, and simplify the link from the sender to the receiver to a noiseless link.
When the noisy channel is a degraded broadcast channel, our model also shares some similarities with source coding with degraded side information (SI) [6, 16]. For some special cases, our model in fact reduces to source coding with degraded SI when an uncoded scheme is used. We compare these two problems with a quadratic Gaussian example. [^1] Throughout this paper, random variables, sample values and their alphabets are denoted by capital, lower and calligraphic letters, e.g., $X,x$ and $\mathcal{X}$. A sequence with length $n$ is denoted by $X^{n}=(X_{1},X_{2},...,X_{n})$. $\mathbb{E}[\cdot]$ is the expectation.
@@
The sender observes the source $S^{n}$ and sends a lossy description to both receivers at rate $R$. The helper observes a correlated source $T^{n}$. It encodes the source into some codeword $X^{l}$ and inputs it to a degraded broadcast channel $P_{YZ|X}=P_{Y|X}P_{Z|Y}$. The receivers observe channel output $Z^{l}$ and $Y^{l}$ and reconstruct $\hat{S}^{n}_{1}$ and $\hat{S}^{n}_{2}$ with distortion levels $D_{1}$ and $D_{2}$, respectively, such that
@@
In some of the results, we restrict the helper to use a separation-based coding scheme. That says, the helper first quantizes the source $T^{n}$ using a lossy source code, and then encodes it using a channel code. The formal definitions of the general joint-source channel coding and separation-based coding schemes are as follows.
@@
  • A pair of decoders $g_{1}:\mathcal{M}\times\mathcal{Z}^{l}\to\hat{S}^{n}_{1},g_{2}:\mathcal{M}\times\mathcal{Y}^{l}\to\hat{S}^{n}_{2}$.
@@
The RDB region $\mathcal{R}_{Se}$ is the set of all $(R,D_{1},D_{2},\rho)$ quadruples that are achievable with a separation-based scheme.
@@
The main result characterizes the tradeoff between the sender’s compression rate, the helper’s broadcast bandwidth, and the reconstruction distortions at two receivers when side information is broadcast over a degraded channel. It shows that a layered (separation-based) scheme is optimal when the helper observes the source directly, yielding simple additive information constraints that fully describe the rate–distortion–bandwidth region.
@@
The outer bound uses a remote source argument that was also used in [17, 7, 20, 12]. The bound is motivated by the rate loss in the source coding with a helper problem. A simple example is the lossless source coding with a helper. The optimal sender-helper rate region is the set of $(R,R_{h})$ s.t.
@@
Let $\mathcal{R}^{in}_{Se}$ be the set of quadruple $(R,D_{1},D_{2},\rho)$ such that there exists a set of finite auxiliary random variables $(U_{1},U_{2},V_{1},V_{2},W)$ such that
@@
The RDB region of the lossy source coding with broadcasted side information under the separation assumption satisfies $\mathcal{R}^{in}_{Se}\subseteq\mathcal{R}_{Se}.$ Furthermore, the cardinalities of the auxiliary random variable alphabets satisfy
@@
The converse is given in Section VI. The achievability follows by constructing lossy descriptions $\{\hat{s}^{n}_{1}\},\{\hat{s}^{n}_{2}\}$ by the marginal distributions $P_{\hat{S}_{1}}P_{\hat{S}_{2}|\hat{S}_{1}}$ and dividing their indices into two parts, sent by the sender and the helper, respectively. The details are provided in Appendix B. In the following, we consider a special case in which the distortion function at the weak receiver is a deterministic distortion function[2][18].
@@
Given a function $\psi:\mathcal{S}\to\widetilde{\mathcal{S}}$ of the source, a distortion function $d:\mathcal{S}\times\hat{\mathcal{S}}\to[0,+\infty)$ is called a deterministic distortion if